area of a regular pentagon inscribed in a circle of radius 12cm.
Answer by khwang(438)

Let AB be a side of the pentagon and O be the center.
then angle AOB = 360/5 = 72 degrees.

And the area of AOB = 1/2 AOsin angle AOB
= 1/2 r^2 sin AOB = 1/2 * 144 sin 72'
= 72 sin 72 = 68.48 cm^2
So the area of the pentagon = 5 area of triangle AOB = 342.38 cm^2
Kenny